Find a polynomial of degree 3 with real coefficients and zeros of -3,-1, and 4, for which f(-2) = - 30.
Answer #1

3 polynomial degree -3,-144 Zerceel, Remo (x-(-3)) (*-(-11) (x-4) f(x) = KC*13) (x+1) (1-4) for No - 2 f(37= 3o - 30 = k (-2+